# On the transitivity of Anosov diffeomorphisms

### v2

@@ Abstract.
We derive a necessary and sufficient condition for an $\Omega-$stable diffeomorphism to be a topologically transitive Anosov diffeomorphism: to exhibit a nonempty intersection of the stable and unstable manifolds of any pair of periodic orbits. To elucidate its dynamical nature, we compare this condition with other properties known to be sufficient for an Anosov diffeomorphism to be topologically transitive. We also describe the $C^{1}$ interior of the set of diffeomorphisms which comply with this condition, discuss examples with a variety of dynamics and present some applications of interest.
@@ 1. Introduction
Let $M$ be a $C^{\infty}$ closed (that is, connected, compact, boundaryless, finite dimensional) Riemannian manifold and $f\colon M\to M$ be a $C^{1}$ diffeomorphism. We say that $f$ is Anosov if $M$ is a hyperbolic set for $f$, meaning that there are constants $c>0$ and $10<\lambda<$ 1 and a $Df-$invariant continuous splitting $T_{x}M=E^{s}_{f}(x)\oplus E^{u}_{f}(x)$ of the tangent space of $M$ at every $x\in M$ so that, for all non-negative integer $n$, one has
@@
A well-known conjecture regarding the classification of Anosov diffeomorphisms states that the non-wandering set of any $C^{1}$ Anosov diffeomorphism is the whole manifold. If true, then any $C^{1}$ Anosov diffeomorphism is topologically transitive and its set of periodic points is dense in the manifold. (We refer the reader to [1] for a more general statement.) And conversely: due to Smale’s spectral decomposition [57], to establish that conjecture it is enough to prove that every Anosov diffeomorphism is topologically transitive. In this work, we provide necessary and sufficient conditions for an Anosov diffeomorphism to be topologically transitive.
Let $(X,d)$ be a compact metric space and $T\colon X\to X$ be a homeomorphism. Given $x\in X$, denote by $\mathcal{O}(x)=\{T^{n}(x)\colon\,n\in\mathbb{Z}\}$ the orbit of $x$ by $T$, and by $\overline{\mathcal{O}(x)}$ its closure in $X$. In what follows, $\Omega(T)$ stands for the set of non-wandering points of $T$, $\operatorname{Per}(T)$ is its subset of periodic points and $\operatorname{Per}_{h}(T)$ denotes its subset of hyperbolic periodic points when $T$ is a diffeomorphism. The recurrent set of $T$ is given by those points in $X$ which are accumulated by their orbits in the future and in the past, that is,
$R(T)=\big{\{}x\in X\colon x\in\alpha(x)\cap\omega(x)\big{\}}.$
The stable and unstable sets of $x\in X$ with respect to $T$ are given respectively by
@@
Clearly, if an Anosov diffeomorphism $f\colon M\to M$ on a smooth closed Riemannian manifold $M$ has the $su-$intersecting property in $M$, then it is topologically transitive. Indeed, by Smale’s spectral decomposition theorem, the set $\Omega(f)$ is a finite union $\Omega_{1}\cup\cdots\cup\Omega_{K}$ of disjoint closed flimit-$f-$invariant pieces, the so called basic sets, on each of which $f$ is topologically transitive. Moreover, each $\Omega_{i}$ contains a dense subset of periodic orbits and, for each $1\leqslant j\leqslant K$, one has $W^{s}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{s}_{f}(x)$, $W^{u}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{u}_{f}(x)$ and $M=\cup_{j=1}^{K}\,W^{s}_{f}(\Omega_{j})=\cup_{j=1}^{K}\,W^{u}_{f}(\Omega_{j})$ (cf. [11, Proposition 3.10]). Take a repeller $\Omega_{i}$ and an attractor $\Omega_{j}$, whose existence is due to the adapted filtration for the spectral decomposition of $\Omega(f)$ (cf. [55, § 2]). It is known that $\Omega_{1}$ contains the stable manifolds of its points and, similarly, $\Omega_{2}$ contains the unstable manifolds of its elements (cf. [11, Lemma 4.9]). So, if $x\in\Omega_{1}$ and $y\in\Omega_{2}$, then $W^{s}_{f}(x)\subset\Omega_{1}$, $W^{u}_{f}(x)\subset\Omega_{2}$ and, by the $su-$intersecting property in $M$, one also has $W^{s}_{f}(x)\cap W^{u}_{f}(y)\neq\emptyset$. Hence $\Omega_{1}\cap\Omega_{2}\neq\emptyset$, which is a contradiction unless $f$ has only one basic set equal to $M$. Thus $f$ is topologically transitive.
@@
It turns out that, if $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$, then $T$ has the barycenter property in $\operatorname{Per}(T)$ (cf. [59][Lemma 2.2]). Moreover, if $X$ has no isolated points and $T$ is expansive, then $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$ if and only if $T$ has the barycenter property in $\operatorname{Per}(T)$. Indeed, the proof of the aforementioned Lemma 2.2 in [59] only requires that for every periodic point $p$ of $T$ there exists $\varepsilon>0$ such that $W^{s}_{T,\,\varepsilon}(p)\subseteq W^{s}_{T}(p)$ and $W^{u}_{T,\,\varepsilon}(p)\subseteq W^{u}_{T}(p)$; and, surely, this condition is valid when $T$ is expansive (cf. [39][Lemma 1]). This ultimately implies that, within the setting of $C^{1}$ diffeomorphisms of a closed Riemannian manifold whose periodic points are hyperbolic, the barycenter property in the set of periodic points is equivalent to the weak $su-$intersecting property in that set.
@@
Our first result concerns homeomorphisms with the shadowing property (see the definition in Subsection 2.7), linking the previous three definitions with the existence of a dense orbit.
@@
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points, and consider $\emptyset\neq A\subseteq X$. Assume that $T$ has the shadowing property.
@@
  • (b)
    If $T(A)=A$, $\Omega(T)\subseteq\overline{A}$ and $T$ has the barycenter property in $A$, then $T$ is topologically transitive.
  • (c)
@@
In particular, if $f\colon M\to M$ is an Anosov diffeomorphism of a closed Riemannian manifold which is either accessible in $R(f)$ or has the barycenter property in $\operatorname{Per}(f)$, then $f$ is topologically transitive.
@@
Assuming that the homeomorphism has the shadowing property and is expansive, we obtain the following correlation between the previous definitions and the specification property (see the definition in Subsection 2.7).
@@
It is known that a $C^{1}$ Anosov diffeomorphism is structurally stable, has the shadowing property, is expansive and its non-wandering set is hyperbolic and coincides with the closure of the set of periodic points (see [5, 11, 49] for more details). We complete this list of properties by deducing the following thorough characterization.
@@
Let $f\colon M\to M$ be an $\Omega-$stable $C^{1}$ diffeomorphism of a smooth closed Riemannian manifold $M$ with dimension at least $2$. Then the following assertions are equivalent:
  • $(a)$
    $f$ has the barycenter property in $\operatorname{Per}(f)$.
  • $(b)$
@@
In [34], M. Lee proved that the $C^{1}$ interior of the diffeomorphisms satisfying the barycenter property in $\operatorname{Per}(f)$ is contained in the set of $\Omega-$stable diffeomorphisms; and that, generically, if a $C^{1}$ diffeomorphism has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega-$stable. We will improve these results by showing that the $C^{1}$ interior of diffeomorphisms satisfying the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$ is equal to the set of topologically transitive Anosov diffeomorphisms (see Proposition 7.10) and by proving that, generically, if a $C^{1}$ diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$ then it is a topologically transitive Anosov diffeomorphism (see Subsection 7.3.2).
Some comments are in order in regard to other known sufficient conditions for an Anosov diffeomorphism to be topologically transitive, and how they compare with the $su-$intersecting property. In [13], M. Brin proved that, if the local product structure of a $C^{1}$ Anosov manifold can be extended to the whole manifold (see the precise meaning of this wording in [13, page 12]), then it is topologically transitive. Actually, Brin’s condition implies that the diffeomorphism has the $su-$intersecting property in the whole manifold, as shown in [13, Remark 4].
Another sufficient condition was proposed by B. Carvalho in [14]: if a $C^{1}$ Anosov diffeomorphism has the two-sided limit shadowing property (see the definition in Subsection 2.7), then it is topologically transitive. Moreover, [14][Lemma 2.4] ensures that if a homeomorphism acting on a compact metric space has the two-sided limit shadowing property, then it has the $su-$intersecting property on the whole space. In Section 3, we present an example of a homeomorphism that has the $su-$intersecting property on the whole domain although it does not satisfy the two-sided limit shadowing property.
In [41, 42], we find conditions of a different nature which also ensure the topological transitivity of an Anosov diffeomorphism. On the one hand, it is known that transitive Anosov diffeomorphisms have a unique measure of maximal entropy (cf. [11]); F. Micena and J. Rodriguez-Hertz proved in [42] that, conversely, if a $C^{1}$ Anosov diffeomorphism has a unique measure of maximal entropy and every point in the manifold is regular in the sense of Oseledets’s theorem [45], then the diffeomorphism is topologically transitive. On the other hand, building upon [11, Theorem 4.14], F. Micena [41] linked the topological transitivity of any $C^{2}$ Anosov diffeomorphism to the Jacobean of the derivative on all the periodic points through the existence of an invariant probability measure equivalent to the volume on the manifold. Micena’s condition is too restrictive, though, since it is known that among $C^{2}$ Anosov diffeomorphisms the ones that admit no invariant probability measure equivalent to the volume on the manifold form an open dense set (cf. [11, Corollary 4.15]).
@@ 2.1. Topological transitivity
A homeomorphism $T$ is topologically transitive if for every nonempty open sets $U,\,V\subset X$ there exists $n\in\mathbb{Z}$ such that $T^{n}(U)\cap V\neq\emptyset$. It is known that this is equivalent to say that $T$ has a dense orbit, that is, there is some $x\in X$ with $\{T^{n}(x)\colon n\in\mathbb{Z}\}$ dense in $X$ (see [4, Theorem 1.4] or [63, Theorem 5.8]).
@@ 2.2. One-sided topologically transitive
A homeomorphism $T$ is one-sided topologically transitive if for every nonempty open sets $U,\,V\subset X$ there exists $n\in\mathbb{N}$ such that $T^{n}(U)\cap V\neq\emptyset$. According to [63, Theorem 5.10], a homeomorphism $T$ of a compact metric space $X$ is one-sided topologically transitive if and only if $T$ is topologically transitive and $\Omega(T)=X$.
@@ 2.4. Chain-transitivity
Given $x,y\in X$, $n\in\mathbb{N}$ and $\delta>0$, a $\delta-$chain of $T$ from $x$ to $y$ of length $n$ is a finite collection $x_{0}=x,\,x_{1},\,\cdots,\,x_{n}=y$ of points in $X$ such that

### v3

# On sufficient conditions for the transitivity of homeomorphisms

@@ Abstract.
We derive a necessary and sufficient condition for a homeomorphism with the shadowing property to be topologically transitive: to have an invariant subset $A$, dense in the non-wandering set, where the barycenter property holds. To elucidate its dynamical nature, we compare this condition with other properties known to be sufficient for an Anosov diffeomorphism to be topologically transitive. We also describe the $C^{1}$ interior of the set of diffeomorphisms which comply with this condition, discuss examples with a variety of dynamics and present some applications of interest.
@@ 1. Introduction
Let $M$ be a $C^{\infty}$ closed (that is, connected, compact, boundaryless, finite dimensional) Riemannian manifold and ${\rm Diff}^{1}(M)$ denote the set of all $C^{1}$ diffeomorphisms of $M$ endowed with the $C^{1}$ topology. We say that $f\in{\rm Diff}^{1}(M)$ is Anosov if $M$ is a hyperbolic set for $f$, meaning that there are constants $c>0$ and $10<\lambda<$ 1 and a $Df-$invariant continuous splitting $T_{x}M=E^{s}_{f}(x)\oplus E^{u}_{f}(x)$ of the tangent space of $M$ at every $x\in M$ so that, for all non-negative integer $n$, one has
@@
A well-known conjecture regarding the classification of Anosov diffeomorphisms states that the non-wandering set of any $C^{1}$ Anosov diffeomorphism is the whole manifold. If true, then any $C^{1}$ Anosov diffeomorphism is topologically transitive and its set of periodic points is dense in the manifold. (We refer the reader to [1] for a more general statement.) And conversely: due to Smale’s spectral decomposition [56], to establish that conjecture it is enough to prove that every Anosov diffeomorphism is topologically transitive. Motivated by a well known sufficient condition for an Anosov diffeomorphism to be topologically transitive, in this work we address other dynamical assumptions on homeomorphisms with the shadowing property under which we also guarantee the existence of a dense orbit.
Let $(X,d)$ be a compact metric space and $T\colon X\to X$ be a homeomorphism. Given $x\in X$, denote by $\mathcal{O}(x)=\{T^{n}(x)\colon\,n\in\mathbb{Z}\}$ the orbit of $x$ by $T$, and by $\overline{\mathcal{O}(x)}$ its closure in $X$. In what follows, $\Omega(T)$ stands for the set of non-wandering points of $T$, $\operatorname{Per}(T)$ is its subset of periodic points and $\operatorname{Per}_{h}(T)$ denotes its subset of hyperbolic periodic points when $T$ is a diffeomorphism. Given $x,y\in X$, $n\in\mathbb{N}$ and $\delta>0$, a $\delta-$chain of $T$ from $x$ to $y$ of length $n$ is a finite collection $x_{0}=x,\,x_{1},\,\cdots,\,x_{n}=y$ of points in $X$ such that
A point $x$ in $X$ is chain-recurrent if, for any $\delta>0$, there exists a $\delta-$chain of $T$ from $x$ to $x$. We denote by $CR(T)$ the set of chain-recurrent points of $T$. The recurrent set of $T$, say $R(T)$, is given by those points in $X$ which are accumulated by their orbits in the future and in the past, that is, $R(T)=\big{\{}x\in X\colon x\in\alpha(x)\cap\omega(x)\big{\}}.$ The stable and unstable sets of $x\in X$ with respect to $T$ are defined respectively by
@@
Clearly, if an Anosov diffeomorphism $f\colon M\to M$ on a smooth closed Riemannian manifold $M$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then it is topologically transitive. Indeed, by Smale’s spectral decomposition theorem, the set $\Omega(f)$ is a finite union $\Omega_{1}\cup\cdots\cup\Omega_{K}$ of disjoint closed flimit-$f-$invariant pieces, the so called basic sets, on each of which $f$ is topologically transitive. Moreover, for each $1\leqslant j\leqslant K$, each $\Omega_{j}$ contains a dense subset of periodic orbits and $W^{s}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{s}_{f}(x)$, $W^{u}_{f}(\Omega_{j})=\cup_{x\,\in\,\Omega_{j}}\,W^{u}_{f}(x)$ and $M=\cup_{j=1}^{K}\,W^{s}_{f}(\Omega_{j})=\cup_{j=1}^{K}\,W^{u}_{f}(\Omega_{j})$ (cf. [11, Proposition 3.10]). Take a repeller $\Omega_{r}$ and an attractor $\Omega_{a}$, whose existence is due to the adapted filtration for the spectral decomposition of $\Omega(f)$ (cf. [55, § 2]). It is known that $\Omega_{r}$ contains the stable manifolds of all its points and, similarly, $\Omega_{a}$ contains the unstable manifolds of its elements (cf. [11, Lemma 4.9]). So, if $x\in\Omega_{r}\cap\operatorname{Per}(f)$ and $y\in\Omega_{a}\cap\operatorname{Per}(f)$, then $W^{s}_{f}(x)\subset\Omega_{r}$, $W^{u}_{f}(x)\subset\Omega_{a}$ and, by the weak $su-$intersecting property in $\operatorname{Per}(f)$, one also has $W^{s}_{f}(x)\cap W^{u}_{f}(y)\neq\emptyset$. Hence $\Omega_{r}\cap\Omega_{a}\neq\emptyset$, which is a contradiction unless $f$ has only one basic set equal to $M$. Thus $f$ is topologically transitive.
@@
It turns out that, if $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$, then $T$ has the barycenter property in $\operatorname{Per}(T)$ (cf. [58, Lemma 2.2]). Moreover, if $X$ has no isolated points and $T$ is expansive, then $T$ has the weak $su-$intersecting property in $\operatorname{Per}(T)$ if and only if $T$ has the barycenter property in $\operatorname{Per}(T)$. Indeed, the proof of the aforementioned Lemma 2.2 in [58] only requires that for every periodic point $p$ of $T$ there exists $\varepsilon>0$ such that $W^{s}_{T,\,\varepsilon}(p)\subseteq W^{s}_{T}(p)$ and $W^{u}_{T,\,\varepsilon}(p)\subseteq W^{u}_{T}(p)$; and, surely, this condition is valid when $T$ is expansive (cf. [39, Lemma 1]). This ultimately implies that, within the setting of $C^{1}$ diffeomorphisms of a closed Riemannian manifold whose periodic points are hyperbolic, the barycenter property in the set of periodic points is equivalent to the weak $su-$intersecting property in that set. For instance, within the family of $C^{1}$ $\Omega-$stable diffeomorphisms we may equally use one or the other property.
@@
Our first result concerns homeomorphisms with the shadowing property (see the definition in Subsection 2.7), linking the three previous definitions with the existence of a dense orbit.
@@
Let $T\colon X\to X$ be a homeomorphism of a compact metric space $(X,d)$ without isolated points, and consider $\emptyset\neq A\subseteq X$.
@@
    If $\overline{A}=X$ and $T$ has the barycenter property in $A$, then $T$ is topologically transitive.
  • (b)
    If $T(A)=A$, $\Omega(T)=\overline{A}$, $T$ has the barycenter property in $A$ and $\Omega(T)=CR(T)$, then $T$ is topologically transitive.
Suppose that $T$ has the shadowing property. Then:
  • (c)
@@
  • (d)
    If $T(A)=A$, $\Omega(T)=\overline{A}$ and $T$ has the barycenter property in $A$, then $T$ is topologically transitive.
  • (e)
@@
We observe that the assumption $\Omega(T)=CR(T)$ holds for homeomorphisms with the shadowing property [6, Theorem 3.1.2]; generically in ${\rm Diff}^{1}(M)$ [10, Corollaire 1.2]; and, by the Closing Lemma [49], for $\Omega$-stable diffeomorphisms if $M$ has dimension $\geqslant 2$. Furthermore, the condition $\Omega(T)=CR(T)$ is not equivalent to shadowing, as exemplified by an irrational rotation on the unit circle, which satisfies $\Omega(T)=CR(T)$ but does not exhibit shadowing.
Every $C^{1}$ Anosov diffeomorphism is structurally stable, has the shadowing property, is expansive, and its non-wandering set is hyperbolic and coincides with the closure of the set of periodic points (cf. [5, 11, 49]). Therefore, we recover from the previous items (b) and (c) the following known result: if $f\colon M\to M$ is an Anosov diffeomorphism of a smooth closed Riemannian manifold $M$ which is either accessible in $R(f)$ or has the barycenter property in $\operatorname{Per}(f)$, then $f$ is a topologically transitive Anosov diffeomorphism (since $\Omega(f)=\overline{\operatorname{Per}(f)}$). Whereas, from item (a) we conclude that, conversely, if $f\colon M\to M$ is a topologically transitive Anosov diffeomorphism, then it has the barycenter property on $M$, since $\Omega(f)=M$ and $f$ has the shadowing property on $M$.
@@
In [34], M. Lee proved that the $C^{1}$ interior of the diffeomorphisms satisfying the barycenter property in $\operatorname{Per}(f)$ is contained in the set of $\Omega-$stable diffeomorphisms; and that, generically, if a $C^{1}$ diffeomorphism has the barycenter property in $\operatorname{Per}(f)$ then it is $\Omega-$stable. We will improve this information by showing that the $C^{1}$ interior of diffeomorphisms satisfying the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$ is equal to the set of topologically transitive Anosov diffeomorphisms (see Proposition 7.10); and by proving that, generically, if a $C^{1}$ diffeomorphism has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$, then it is a topologically transitive Anosov diffeomorphism (see Subsection 7.3.2).
Assuming the shadowing property and expansivity, we obtain the following correlation between the previous definitions and the specification property (see the definition in Subsection 2.7).
@@
For any $C^{1}$ $\Omega-$stable diffeomorphism, the next result establishes a general equivalence between the topological transitivity and several other dynamical properties.
@@
Let $M$ be a smooth closed Riemannian manifold, with dimension at least $2$, and $f\colon M\to M$ be a $C^{1}$ $\Omega-$stable diffeomorphism. The following statements are equivalent:
  • $(1)$
    $f$ has the two-sided limit shadowing property.
  • $(2)$
    $f$ has the $su-$intersecting property in $M$.
  • $(3)$
    $f$ has the $su-$intersecting property in $\operatorname{Per}(f)$.
  • $(4)$
    $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$.
  • $(5)$
    There exists $A\subseteq M$ such that $f(A)=A$, $\Omega(f)=\overline{A}$ and $f$ has the barycenter property in $A$.
  • $(6)$
@@
  • $(7)$
    $f$ has the specification property.
  • $(8)$
    $f$ has the almost specification property.
  • $(9)$
    $f$ has the asymptotic average shadowing property.
  • $(10)$
    $f$ has the average shadowing property.
  • $(11)$
    $f$ is chain-transitive.
  • $(12)$
    $\mathcal{P}_{f}^{erg}(M)$ is dense in $\mathcal{P}_{f}(M)$.
  • $(13)$
    $f$ has the gluing orbit property.
Some comments are in order regarding other known sufficient conditions for an Anosov diffeomorphism to be topologically transitive, and how they compare with the $su-$intersecting property. In [13], M. Brin proved that, if the local product structure of a $C^{1}$ Anosov manifold can be extended to the whole manifold (see the precise meaning of this wording in [13, page 12]), then it is topologically transitive. Actually, Brin’s condition implies that the diffeomorphism has the $su-$intersecting property in the whole manifold, as shown in [13, Remark 4]. Another sufficient condition was proposed by B. Carvalho in [14]: if a $C^{1}$ Anosov diffeomorphism has the two-sided limit shadowing property (see the definition in Subsection 2.7), then it is topologically transitive. Moreover, [14, Lemma 2.4] ensures that if a homeomorphism acting on a compact metric space has the two-sided limit shadowing property, then it has the $su-$intersecting property on the whole space. In Section 3, we present an example of a homeomorphism that has the $su-$intersecting property on the whole domain although it does not satisfy the two-sided limit shadowing property. In [41, 42], we find conditions of a different nature which also ensure the topological transitivity of an Anosov diffeomorphism. On the one hand, it is known that transitive Anosov diffeomorphisms have a unique measure of maximal entropy (cf. [11]); F. Micena and J. Rodriguez-Hertz proved in [42] that, conversely, if a $C^{1}$ Anosov diffeomorphism has a unique measure of maximal entropy and every point in the manifold is regular in the sense of Oseledets’s theorem [45], then the diffeomorphism is topologically transitive. On the other hand, building upon [11, Theorem 4.14], F. Micena [41] linked the topological transitivity of any $C^{2}$ Anosov diffeomorphism to the Jacobean of the derivative on all the periodic points through the existence of an invariant probability measure equivalent to the volume on the manifold. Micena’s condition is too restrictive, though, since it is known that among $C^{2}$ Anosov diffeomorphisms the ones that admit no invariant probability measure equivalent to the volume on the manifold form an open dense set (cf. [11, Corollary 4.15]).
@@ 2.1. Topological transitivity
A homeomorphism $T$ is topologically transitive if for every nonempty open sets $U,\,V\subset X$ there exists $n\in\mathbb{Z}$ such that $T^{n}(U)\cap V\neq\emptyset$. It is known that this is equivalent to say that $T$ has a dense orbit, that is, there is some $x\in X$ with $\{T^{n}(x)\colon n\in\mathbb{Z}\}$ dense in $X$ (see [4, Theorem 1.4] or [62, Theorem 5.8]).
@@ 2.2. One-sided topologically transitive
A homeomorphism $T$ is one-sided topologically transitive if for every nonempty open sets $U,\,V\subset X$ there exists $n\in\mathbb{N}$ such that $T^{n}(U)\cap V\neq\emptyset$. According to [62, Theorem 5.10], a homeomorphism $T$ of a compact metric space $X$ is one-sided topologically transitive if and only if $T$ is topologically transitive and $\Omega(T)=X$.